Hi all, Im having a bit of trouble with Matrix problems at the moment.
Can someone please help me out with the following?
Write this system of equations in matrix form
y + z = 0 ,
−x + y − z = 9 ,
x − z = 4 .
Iwant to try and get a better understanding of the solving in matrix form, so steps in working nd explanations would be really helpful.
Thanks. -Nick

Answer by anantha(86)

sol:
matrix equations obtained from the equations
a1x+b1y+c1z+d1=0,a2x+b2y+c2z+d2=0,a3x+b3y+c3z+d3=0
is [a1 b1 c1] [x ] [ -d1]
[a2 b2 c2] [y ] = [ -d2 ]
[a3 b3 c3] [z ] [ -d3 ]
system of equations in this problem
0x+y+z=0
-x+y-z=9
x+0y-z=4
comparing these three equations as above equations
a1x+b1y+c1z+d1=0
0x +1y+1z=0
here a1=0,b1=1,c1=1,d1=0
a2x+b2y+c2z+d2=0
-1x+1y-1z-9=0
here a2=-1,b2=1,c2=-1,d2=-9
a3x+b3y+c3z+d3=0
1x+0y-1z-4=0
here a3=1,b3=0,c3=-1,d3=-4
substitute these values in the matrix form
[0 1 1] [x] [0]
[-1 1 -1] [y] = [9]
[1 0 -1] [z] [4]